OpenStudy (anonymous):

find the slope and y-intercept 2x-y=5 n graph line

OpenStudy (anonymous):

if you arrange it like y=2x-5 than you can read off the slope. It is the multiplier of x, so it is 2. For the y-intercept, you have to set y=0 and than 2x-0=5 so x=5/2

OpenStudy (amistre64):

-2/-1 = slope = 2 yint = x=0; -5

OpenStudy (anonymous):

oh yeah my was the x intercept

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

the slope is; ax+by = c -a/b

OpenStudy (amistre64):

equation of a line is equation of a line :)

OpenStudy (anonymous):

so how do i wite the equation

OpenStudy (amistre64):

y = (-a/b)x +c slope = -a/b

OpenStudy (amistre64):

depends on which form of the equation you want to write

OpenStudy (amistre64):

ax +by - c = 0 ax + by = c y = -(a/b)x + c y-y1 = (-a/b)(x-x1)

OpenStudy (amistre64):

x = x1 + at y = y1 + bt z = 0

OpenStudy (anonymous):

how did you find the y intercept

OpenStudy (amistre64):

what is the definition of a y intercept?

OpenStudy (amistre64):

what value is y when the graph is caught by the y axis ? the y axis is the x=0 axis what is the value of the equation when x=0?

OpenStudy (amistre64):

2(0) - y = 5 -y = 5 y = -5

OpenStudy (amistre64):

(0,-5)

OpenStudy (anonymous):

oh ok so what was the -2/-1 for